OpenStudy (anonymous):

What is the value of y when x = -3? 4x + 2y = 8 is it 2

OpenStudy (anonymous):

when x = -3, 4(-3) + 2y = 8 -12 + 2y = 8 2y = 20 y = 10 it cannot be 2

OpenStudy (anonymous):

i subtracted and didnt add simple mistake i actually got -2 but i understand now

OpenStudy (anonymous):

ok then as long as you understand

OpenStudy (anonymous):